• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M.Iodice

    Log entry time 07:56:48 on June16,2005

    Entry number 146218

    keyword=Owl shift summary

    00:30 - end of control access to fix the vacuum valves (by an MCC tech)
    01:20 - MCC tried to send CW beam without success and turned to
    tune beam
    01:40 - CW beam appears
    01:40 - 01:45 - Nilanga takes data with empty target
    01:51 - Waterfall target, 50uA, running at different pump speeds with the HAPPEX DAQ and the HRS DAQ
    02:42 - end of the target thickness tests
    02:46 - 03:10 - Nilanga's tests with the thin carbon target
    03:36 - During the last hour we saw only 10 min of beam
    MCC running a pulsed beam to Hall C...
    03:44 - beam appears. Started tuning for Moller measurements
    06:20 - finished with Moller measurements
    We asked for full survey and restricted access.
    07:30 - Sweep completed